How to find range of sinx-cosx

Grade:11

4 Answers

radhika
8 years ago
asinx – bcosx
min value = -√a2+b= -√1+1 = -√2
max value = √a2+b2 = √2

 
AUREA
8 years ago
f(x) = sinx – cosx
     =√2 x 1/√2 ( sinx – cosx)
     =√2 x { (1/√2)sinx – (1/√2)cosx }
     =√2 sin (x – pi/4)
range of sine function is [-1,1]
so range of f(x) will be [-√2,√2]
grenade
8 years ago
The max value equals root2 and minimum minus root 2
Thence the range is between min and maz
Hence the answer would be from minus root 2to root 2
ng29
8 years ago
just find the max and min values of this equation by differentiating it 
we get min = -(2)1/2  and max =(2)1/2
so its range of function 
 
approve if useful